mark is five times older than his sister.if the sum of their ages three years ago is twenty seven. how old is mark
Answer by Menjax(62)

Let m be marks current age
Let s be his sisters current age
m=5s #1
(m-3) + (s-3) = 27 #2
From 1
5s-3+s-3=27
6s=33
s=11/2
m=55/2